OpenStudy (anonymous):

A line has a slope of and a y-intercept of 3. Which equations represent the line? Choose all answers that are correct. need help asap got to go soon plz help mee!!!:) fan and metal if helped

OpenStudy (anonymous):

what is the slope?

OpenStudy (anonymous):

-1/2

OpenStudy (anonymous):

can you see the answer choices

OpenStudy (anonymous):

nope the link leads to a google page with nothing but from what you have given i think it should be ok using y=mx+b equation would be \[y=-\frac{ 1 }{ 3 }x+3\]

OpenStudy (anonymous):

is it any of the options provided

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

sorry typing error it is\[y=-\frac{ 1 }{ 2 }x+3\]

OpenStudy (anonymous):

a) 2y=6+x b) x+2y=6 c)y=-1/2x+3 d)(y-2)=-1/2(x-2)

OpenStudy (anonymous):

ok c is the right option

OpenStudy (anonymous):

ok any more it is a chose all that are correct

OpenStudy (anonymous):

b is also right if you rearrange

OpenStudy (anonymous):

ok thanks can you help me with more plz i will open new thred

OpenStudy (anonymous):

I really need to go maybe next time oh and d is also right

OpenStudy (anonymous):

ok thanks it is ok